Circular No. 6796

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 971227 A. Coletta, G. Gandolfi, and M. Smith, BeppoSAX Science Operation

Center, Telespazio, Rome; L. Piro, M. Cinti, and P. Soffitta, Istituto di Astrofisica Spaziale, CNR, Rome; and J. Heise, Space Research Organisation, Utrecht, report on behalf of the BeppoSAX team: “A γ-ray burst (Bacodine trigger 6546) has been detected by the BeppoSAX Gamma-Ray burst Monitor on Dec. 27.34938 UT. A quick-look analysis shows a structured γ-ray burst lasting ∼ 10 s with a peak flux of 630 counts/s. The γ-ray burst has been detected also by the Wide Field Camera WFC2 with a peak flux of ∼ 1.8 Crab. The position given by the quick-look analysis is α = 12^{h}57^{m}29^{s}, δ = +59^{o}16′.3 (equinox 2000.0). The event was detected in a part of the orbit when the satellite was not in an optimal pointing configuration, resulting in an estimated error radius of 10′. A BeppoSAX Narrow Field follow-up observation is in progress. Observations at all wavelengths are urged.”

SUPERNOVA 1997ei IN NGC 3963 P. Garnavich, S. Jha, R. Kirshner, and P. Challis, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum obtained by L. Macri with the 1.5-m Tillinghast telescope on Dec. 27.5 UT shows that SN 1997ei is a type-Ia supernova, ∼ 1 week after maximum. Narrow emission lines from the host galaxy provide a redshift of z = 0.011.

GRB 971214 N. Tanvir, Cambridge University; R. Wyse, Johns Hopkins University;

G. Gilmore, Cambridge University; C. Corson, Apache Point Observatory; and the APO GRB consortium write: “We obtained the following magnitude of GRB 971214 with the APO 3.5-m telescope: Dec. 15.44 UT, J = 20.27 ± 0.25. Combined with the APO R-band measurement (IAUC 6791) on Dec. 15.51, this implies a color of R − J = +1.8 ± 0.3. Observations by the same team on Dec. 16.45 provide a tentative detection at J = 21.5.”

COMET 55P/TEMPEL-TUTTLE Visual m_{1} and coma-diameter estimates: Dec. 26.81 UT, [12.0, – (A.

Pearce, Cadoux, W. Australia, 0.41-m reflector); 27.48, 11.0, 4′ (A. Hale, Cloudcroft, NM, 0.41-m reflector); 27.79, 12.2, 3′.0 (S. Yoshida, Ibaraki, Japan, 0.46-m reflector); 27.81, 12.1, 3′ (Pearce); 28.81, 11.9, 3′.6 (Pearce).

1997 December 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green